You are tasked with evaluating the purchase of a vending machine for the snack room.

The base price is $4,000 and it would cost another $1,000 to modify the machine to install the machine. The equipment falls in the MACRS 3 year class of depreciation with rates of 33%, 45%, 15% and 7%. The machine would require an investment in snacks (inventory/net operating working capital) of $500. The machine would produce revenue of $3,000 per year with costs of $534 per year. The firm’s tax rate is 20%.

The firm plans on selling the machine at the end of three years for $1500 and recovering their investment in net working capital at that time as well.

What is the operating cash flow associated with the first year of operation?
Answer #1
operating cash flow associated with the first year of operation
i revenue 3000
ii Cost 534
iii Depreciation =(4000+1000)*33% 1650
iv=i-ii-iii Profit before tax 816
v=iv*20% tax @ 20% 163.2
vi=iv-v Profit after tax 652.8
vii=vi+iii Operating cash flow 2302.8
Therefore operating cash flow 2302.8